600 ml of a mixture of O3 and O2 weighs 1g at NTP. Calculate the volume of ozone in the mixture. Please give solution.

600 ml of  a mixture of O3  and 
O2  weighs 1g at NTP. Calculate 
the volume of ozone in the mixture. Please give solution.

Grade:11

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Solution:
Let us assume x g of oxygen and y g of ozone.
Hence the moles for oxygen and ozone would be equal to:
\frac{x}{32}; \frac{y}{48}
Now the volume occupied by oxygen and ozone would be equal to:
\frac{x}{32}\times 22.4; \frac{y}{48}\times 22.4
Therefore,
we get:x=0.57 and y=0.428
So the volume of ozone would be equal to:0.2 litres.
Thanks & Regards
Sumit Majumdar,
askIITians Faculty
Ph.D,IIT Delhi